# Prompt Engineer: Analyzing Skill Requirements in the AI Job Market

### v1

@@ 4.3.3 Correlation between soft and hard skills of prompt engineers
We identified two significant negative co-occurrences of skills. The skills Teamwork & Collaboration are negatively correlated with SQL & Database Systems ($r=-2.23$). This may imply that prompt engineers working on back-end–focused and database-oriented tasks are less integrated with team processes. The second negative association is between Stakeholder Management & Business Acumen and Machine Learning & AI ($r=-2.53$). This could indicate that prompt engineers prioritize business-facing tasks over heavy technical AI work. It could, however, also indicate a potential technical sub-role within the prompt engineer role, where machine learning and developer-oriented tasks are emphasized over higher-level communication and business tasks.
Figure 8: Skills co-occurrence between soft and hard skills for prompt engineer positions. High values indicate pairs of skills that are often mentioned together in job descriptions, whereas low values indicate the opposite. Values of $|r|>2.0$ indicate significant deviations from expected values ($p<0.05$).
@@ 5 Discussion
In this section, we discuss our findings on the emergence of prompt engineering as a profession, its skill requirements, and implications for the job market.
@@ 5.1 The Emerging Status of Prompt Engineering
Our analysis shows that prompt engineering remains in its early stages as an emerging profession. With just 72 prompt engineer positions identified (less than 0.5% of the related roles in our dataset), the role has not yet achieved widespread adoption across industries. This limited presence confirms that despite significant industry attention on large language models, dedicated prompt engineering positions are still relatively rare. This finding aligns with the timeline of LLM development and adoption. While LLMs have gained rapid attention since late 2022, the establishment of specialized roles typically lags behind technology adoption as organizations first experiment with assigning these new responsibilities to existing staff before creating dedicated positions.
@@ 5.2 The Distinct Soft Skill Profile of Prompt Engineers
While skills may become obsolete (Horton and Tambe, 2025), new skills can emerge, giving rise to professions with a distinct skill profile. This skill profile has a strong overlap with what has been termed the “21st-century” skill profile in prior literature (van Laar et al., 2017; Federiakin et al., 2024). For instance, van Laar et al. (2017) identified seven 21st-century skills in their systematic literature review: technical, information management, communication, collaboration, creativity, critical thinking, and problem solving. This set of skills is well represented in our analysis of the prompt engineering profession. The two key soft skills identified for the emerging role of prompt engineers are an enhanced focus on communication (21.9%) and creative problem-solving (15.8%). Prompt engineer positions emphasize strong communication skills more than the other investigated job roles. This reflects the expectation that prompt engineers must effectively communicate their innovative approaches and solutions to various stakeholders. Prompt engineers are also expected to apply creative approaches to problem-solving, with skills such as “creative mindset” and “creative problem-solving” appearing frequently in job postings. This suggests employers recognize that prompting is an innovative field that requires engineers to think creatively when solving problems.
@@
Our analysis reveals that while communication and teamwork capabilities are universal soft skills in tech roles—aligning with Lievens and Sackett (2012)’s findings on interpersonal skills—prompt engineers show distinctively high proportions of these requirements. As 6(a) illustrates, Communication Skills and Teamwork & Collaboration each represent 21.9% of soft skill requirements, with Problem-solving Abilities (15.8%) also emphasized more strongly compared to other job roles. Interestingly, Time & Project Management skills appear least frequently (1.9%), which may align with prompt engineering being an innovative emerging field that lacks standardized management processes. These findings indicate that prompt engineers serve as innovation communicators at the interface between LLM capabilities and stakeholder needs.
@@
Our analysis also shows that the skills profile of prompt engineers is only partially reflected in existing skills taxonomies, such as the UK skills taxonomy (Djumalieva and Sleeman, 2018) and the Burning Glass skills taxonomy (Burning Glass Technologies, 2019). Prompt engineers have a specific profile that not only leverages advanced linguistic expertise—including nuanced control of semantics, syntax, and discourse strategies—to iteratively craft and optimize prompts that effectively steer AI behavior, a dimension largely absent from traditional skills frameworks. In addition, prompt engineers, use tools and frameworks with scripting languages, as demonstrated by our analysis of hard skills. In the following section, we expand our discussion on the technical skills profile of prompt engineers.
@@ 5.3 Technical Skills Defining Prompt Engineering
The analysis of hard skills provides strong evidence that prompt engineering is emerging as a distinct technical discipline rather than simply an extension of existing job roles. Chi-square testing revealed highly significant differences among the skill profiles of different job types ($\chi^{2}(44,N=20662)=67291.55$, $p<0.0001$), confirming that these technical skill profiles represent meaningful distinctions. Notably, the side-by-side comparison with other job roles (cf. Table 4, Table 4, and 7(b)) highlights that prompt engineers stand out in four technical categories:
@@
    Agile & Testing (18.7%): Prompt engineers need significant more testing skills than other roles. Data analysts, for instance, only require this skill in 8.5% of the job postings. This matches the findings by Liu and Chilton (2022) – the prompt engineering process involves extensive testing and optimization.
@@
In addition, prompt engineers require Python & Scripting (9.2%) skills. While Python is often used for ETL, statistical analysis, machine learning pipelines, or data visualization in traditional job roles, prompt engineers use Python specifically to interface with and manipulate large language models (LLMs). For instance, prompt engineers may use Python to orchestrate API calls to LLMs, construct dynamic prompts, parse model outputs, and evaluate generations – tasks that are more about language manipulation and real-time interaction than batch data processing. Prompt engineers also write lightweight scripts to prototype interactions, automate prompt variation, or integrate LLMs into applications, often with Python-based tools and libraries, such as requests, langchain, or openai.
@@
This unique mix of skills shows prompt engineering is becoming a specialized field focused on making language models work well in real applications.
@@ 5.4 Implications for the Job Market
The emergence of prompt engineering as a distinct profession with specialized skill requirements has several implications for both job seekers and employers:
@@
    Hybrid skill set requirements: Prompt engineers need a unique combination of technical AI knowledge, strong communication, and strong creative skills. This hybrid profile is rare, as it blends competencies from traditionally separate domains—such as software engineering, linguistics, and design. In practice, prompt engineers must be capable of understanding model behavior, crafting precise language instructions, and iteratively refining outputs based on feedback. For employers, this means recruiting for a complex and still evolving role that few candidates are formally trained for, making talent acquisition and role definition more difficult.
@@ 5.5 Limitations and Future Work
    Sample size: With only 72 prompt engineer positions identified on LinkedIn, our analysis gives insights into the emerging landscape of prompt engineering in an important online job market, but the result may not be indicative of wider developments.
@@ 6 Conclusion
Our data-driven analysis of prompt engineering job postings reveals the role of prompt engineer remains rare, comprising less than 0.5% of related positions despite growing LLM adoption. The emerging profession shows a distinctive skill profile combining strong communication and creative problem-solving with specialized AI knowledge. Most surprisingly, prompt engineering extends far beyond crafting textual prompts. The emphasis on testing and DevOps skills indicates that prompt engineers are professionals responsible for the entire solution lifecycle—designing, testing, integrating, and deploying LLM systems into production workflows. This creates both opportunities and challenges, with important caveats about the job role’s future. The combination of slow organizational adoption and rapid AI evolution creates uncertainty about the longevity of dedicated prompt engineering positions. As LLMs improve and no-code interfaces advance, some current prompt engineering tasks may become automated. Job seekers should develop broader AI foundations alongside prompt engineering skills, while employers may favor upskilling existing employees over creating specialized roles.
@@
Appendix
Appendix A List of Countries
United States (6553; 31.7%), India (3747; 18.1%), United Kingdom (1490; 7.2%), Canada (621; 3%), Singapore (514; 2.5%), Poland (494; 2.4%), Portugal (440; 2.1%), Spain (439; 2.1%), Germany (434; 2.1%), Philippines (418; 2%), Australia (347; 1.7%), Mexico (276; 1.3%), Israel (273; 1.3%), Netherlands (269; 1.3%), Romania (210; 1%), Brazil (209; 1%), France (199; 1%), Malaysia (195; 0.9%), Belgium (183; 0.9%), Ireland (176; 0.9%), Indonesia (160; 0.8%), China (152; 0.7%), Vietnam (141; 0.7%), Sweden (140; 0.7%), Colombia (136; 0.7%), South Africa (134; 0.6%), Italy (127; 0.6%), Hungary (122; 0.6%), Argentina (122; 0.6%), Egypt (117; 0.6%), Greece (114; 0.6%), Japan (109; 0.5%), Thailand (102; 0.5%), Ukraine (98; 0.5%), Czech Republic (97; 0.5%), Turkey (96; 0.5%), United Arab Emirates (93; 0.5%), Lithuania (89; 0.4%), Bulgaria (88; 0.4%), Switzerland (88; 0.4%), Taiwan (87; 0.4%), Pakistan (72; 0.3%), Saudi Arabia (70; 0.3%), Denmark (58; 0.3%), Finland (49; 0.2%), Serbia (47; 0.2%), Cyprus (45; 0.2%), Slovakia (39; 0.2%), New Zealand (38; 0.2%), Morocco (34; 0.2%), Nigeria (29; 0.1%), Luxembourg (29; 0.1%), South Korea (27; 0.1%), Sri Lanka (27; 0.1%), Estonia (26; 0.1%), Austria (26; 0.1%), Malta (22; 0.1%), Norway (17; 0.1%), Kenya (17; 0.1%), Croatia (17; 0.1%), Latvia (14; 0.1%), Bangladesh (14; 0.1%), Chile (13; 0.1%), Slovenia (11; 0.1%), Qatar (11; 0.1%), Tunisia (6; $<$0.1%), Russia (2; $<$0.1%), Belarus (2; $<$0.1%), Liechtenstein (1; $<$0.1%)

### v2

@@ 4.3.3 Correlation between soft and hard skills of prompt engineers
We identified two significant negative co-occurrences of skills. The skills Teamwork & Collaboration are negatively correlated with SQL & Database Systems (r=2.23r=-2.23). This may imply that prompt engineers working on back-end–focused and database-oriented tasks are less integrated with team processes. The second negative association is between Stakeholder Management & Business Acumen and Machine Learning & AI (r=2.53r=-2.53). This could indicate that prompt engineers prioritize business-facing tasks over heavy technical AI work. It could, however, also indicate a potential technical sub-role within the prompt engineer role, where machine learning and developer-oriented tasks are emphasized over higher-level communication and business tasks.
Figure 5: Skills co-occurrence between soft and hard skills for prompt engineer positions. High values indicate pairs of skills that are often mentioned together in job descriptions, whereas low values indicate the opposite. Values of $|r|>2.0$ indicate significant deviations from expected values ($p<0.05$). Source: Authors own work.
@@ 5 Discussion
We humans increasingly rely on generative AI, and in particular conversational AI. In our daily, but also our professional interaction with large language models, we employ “prompt engineering” almost routinely to make the LLM produce our intended outcomes. The practice and skill of prompt engineering (Oppenlaender et al., 2024) is playing an increasingly important role in research and industry. In research, some have cautioned against the growing “LLMification” that is taking place in some research areas (Pang et al., 2025). In industry, pre-trained LLMs have unlocked capabilities and opportunities that were previously not feasible or cost-effective. Without a doubt, technological change (World Economic Forum, 2025) and our increased acceptance and reliance on AI will contribute and accelerate changes in skills and jobs. The World Economic Forum’s 2025 report on the future of jobs (World Economic Forum, 2025) highlights the close linkage of these skills with the digital transformation. There will be changes in skills in the work force in the coming years, with new job roles being created (Capps, 2025). In this work, we examined a snapshot of the current state of prompt engineering in industry, and took inventory of the skills required by prompt engineers today. In the remainder of this section, we discuss our findings on the emergence of prompt engineering as a profession in industry, its skill requirements, and implications for the job market.
@@ 5.1 The Emerging Status of Prompt Engineering
Our analysis shows that prompt engineering remains in its early stages as its own profession. With only 72 prompt engineer positions identified (less than 0.5% of the related roles in our dataset), the role has not yet achieved widespread adoption across industries. This limited presence confirms that despite significant industry attention on large language models, dedicated prompt engineering positions are still relatively rare. This finding aligns with the timeline of LLM development and adoption. While LLMs have gained rapid attention since late 2022, the establishment of specialized roles typically lags behind technology adoption as organizations first experiment with assigning these new responsibilities to existing staff before creating dedicated positions.
@@ 5.2 The Distinct Soft Skill Profile of Prompt Engineers
While technical skills may become obsolete (Horton and Tambe, 2025), new skills can emerge, giving rise to professions with a distinct skill profile. The skill profile of prompt engineers has a strong overlap with what has been termed the “21st-century” skill profile in prior literature (van Laar et al., 2017; Federiakin et al., 2024). For instance, van Laar et al. identified seven 21st-century skills in their systematic literature review (van Laar et al., 2017): technical, information management, communication, collaboration, creativity, critical thinking, and problem solving. In their report on the Future of Jobs, the World Economic Forum referred to these skills as “Core skills in 2030” (World Economic Forum, 2025, p. 41), including skills such as creative thinking, technological literacy, and AI and big data. We find these skills are already well represented in our analysis of prompt engineering positions today.
The two key soft skills identified for the emerging role of prompt engineers are an enhanced focus on communication (21.9%) and creative problem-solving (15.8%). Prompt engineer positions emphasize strong communication skills more than the other investigated job roles. This reflects the expectation that prompt engineers must effectively communicate their innovative approaches and solutions to various stakeholders. Prompt engineers are also expected to apply creative approaches to problem-solving, with skills such as “creative mindset” and “creative problem-solving” appearing frequently in job postings. This suggests employers recognize that prompting is an innovative field that requires engineers to think creatively when solving problems.
@@
Our analysis reveals that while communication and teamwork capabilities are universal soft skills in tech roles, aligning with Lievens and Sackett’s findings on interpersonal skills (Lievens and Sackett, 2012), prompt engineers show distinctively high proportions of these requirements. As Figure 3a illustrates, Communication Skills and Teamwork & Collaboration each represent 21.9% of soft skill requirements, with Problem-solving Abilities (15.8%) also emphasized more strongly compared to other job roles. Interestingly, Time & Project Management skills appear least frequently (1.9%), which may align with prompt engineering being an innovative emerging field that lacks standardized management processes. These findings indicate that prompt engineers serve as innovation communicators at the interface between LLM capabilities and stakeholder needs.
@@
Our analysis also shows that the skills profile of prompt engineers is only partially reflected in existing skills taxonomies, such as the UK skills taxonomy (Djumalieva and Sleeman, 2018) and the Burning Glass skills taxonomy (Burning Glass Technologies, 2019). Prompt engineers have a specific profile that not only leverages advanced linguistic expertise, including nuanced control of semantics, syntax, and discourse strategies, to iteratively craft and optimize prompts that effectively steer AI behavior, a dimension largely absent from traditional skills frameworks. In addition, prompt engineers use tools, frameworks, and scripting languages, as demonstrated by our analysis of hard skills. In the following section, we expand our discussion on the technical skills profile of prompt engineers.
@@ 5.3 Technical Skills Defining Prompt Engineering
The analysis of hard skills provides strong evidence that prompt engineering is technically distinct, rather than simply an extension of existing job roles. Chi-square testing revealed highly significant differences among the skill profiles of different job types ($\chi^{2}(44,N=20662)=67291.55$, $p<0.0001$), confirming that these technical skill profiles represent meaningful distinctions. Notably, the side-by-side comparison with other job roles (cf. Table 4, Table 4, and Figure 4b) highlights that prompt engineers stand out in four technical categories:
@@
    Agile & Testing (18.7%): Prompt engineers need significant more testing skills than other roles. Data analysts, for instance, only require this skill in 8.5% of the job postings. This matches the findings by Liu and Chilton (Liu and Chilton, 2022) – the prompt engineering process involves extensive testing and optimization.
@@
In addition, prompt engineers require Python & Scripting (9.2%) skills. While Python is often used for ETL, statistical analysis, machine learning pipelines, or data visualization in traditional job roles, prompt engineers use Python specifically to interface with and manipulate large language models (LLMs). For instance, prompt engineers may use Python to orchestrate API calls to LLMs, construct dynamic prompts, parse model outputs, and evaluate generations – tasks that are more about language manipulation and real-time interaction than batch data processing. Prompt engineers also write lightweight scripts to prototype interactions, automate prompt variation, or integrate LLMs into applications, often with Python-based tools and libraries, such as LangChain or OpenAI.
@@
This unique mix of skills shows prompt engineering is focused on making language models work well in real applications, with emerging consequences for the job market.
@@ 5.4 Implications for the Job Market
The emergence of prompt engineering as a distinct and specialized skill profile has several implications for both job seekers and employers:
@@
    Hybrid skill set requirements: Prompt engineers need a unique combination of technical AI knowledge, strong communication, and strong creative skills. This hybrid profile is rare, as it blends competencies from traditionally separate domains, such as software engineering, linguistics, and design. In practice, prompt engineers must be capable of understanding model behavior, crafting precise language instructions, and iteratively refining outputs based on feedback. For employers, this means recruiting for a complex and still evolving role that few candidates are formally trained for, making talent acquisition and role definition more difficult.
@@ 5.5 Limitations and Future Work
    Sample size: With only 72 prompt engineer positions identified on LinkedIn at our snapshot date, our analysis gives insights into the emerging landscape of prompt engineering in an important online job market, but the result may not be indicative of wider developments.
@@ 6 Conclusion
Our data-driven analysis of prompt engineering job postings reveals the role of prompt engineer remains rare, comprising less than 0.5% of related positions despite growing LLM adoption. Prompt engineering roles show a distinctive skill profile combining strong communication and creative problem-solving with specialized AI knowledge. Most surprisingly, prompt engineering extends far beyond crafting textual prompts. The emphasis on testing and DevOps skills indicates that prompt engineers are professionals responsible for the entire solution lifecycle: designing, testing, integrating, and deploying LLM systems into production workflows. This creates both opportunities and challenges, with important caveats about the job role’s future. The combination of slow organizational adoption and rapid AI evolution creates uncertainty about the longevity of dedicated prompt engineering positions. As LLMs improve and no-code interfaces advance, some current prompt engineering tasks may become automated. Job seekers should develop broader AI foundations alongside prompt engineering skills, while employers may favor upskilling existing employees over creating specialized roles.
@@
Disclosure of interest
The authors report there are no competing interests to declare.